use implicit differentiantion to find the derivative of the fuction

cos(x^2+2y) + xe^y = 0


1
Expert's answer
2020-10-08T16:04:39-0400

"cos(x^2+2y) + xe^y = 0"


"\frac{d}{dx}[cos(x^2+2y) + xe^y]=\frac{d}{dx}[0]"


"\frac{d}{dx}[cos(x^2+2y) + xe^y]=\frac{d}{dx}[0]"


"(-sin(x^2+2y)).\frac{d}{dx}[x^2+2y]+\frac{d}{dx}[e^y].x+e^y.\frac{d}{dx}[x]=0"


"-(\frac{d}{dx}[x^2]+\frac{d}{dx}[2y])sin(x^2+2y)+e^y.\frac{d}{dx}[y].x+e^y.1=0"


"-(2x+2y\prime)sin(x^2+2y)+e^yy\prime x+e^y=0"


"-(2y\prime +2x)sin(2y+x^2)+xe^yy\prime +e^y=0"


"y\prime =\frac{2xsin(2y+x^2)-e^y}{2sin(2y+x^2)-xe^y}"